# Revisiting the Galactic Winds in M82 I: the recent starburst and launch of outflow in simulations

### v2

@@ 2.4.4 Core-collapse SNe
Conventionally, stars within the zero-age main sequence mass range of $8\sim 30M_{\odot}$ are thought to end their life through the outbreak of core-collapse supernovae (CCSN). Most previous models use a fixed thermal energy injection value of $10^{51}$ erg. However, recent studies indicate that the energy and mass of the ejecta as well as the composition of the ejecta depend on certain parameters, especially the ZAMS mass of the progenitor. In our study, we mainly follow the work of Gutcke et al. (2021). We divide the total mass of each star particle (star cluster) into multiple ZAMS stars according to a Kroupa (2002) IMF. In practice, instead of creating many refined small particles, we store a list of individual virtual stars’ information, including mass and expected lifetime, for each star particle (star cluster). At each time step, we update the age of virtual stars and loop over the lists to locate all the massive stars that have exceeded their expected lifetime and then trigger either a CCSN or failed SN for each such virtual star. The injected energy, ejecta mass, and metal of each SN (failed SN) are then calculated from the Sukhbold et al. (2015) CCSN model accordingly. Stellar wind and UV radiation before the outbreak of CCSN have already cleared the path for SNe Dale et al. (2014). Therefore, we directly inject 100% of SNe’s mass, metal, and energy output, together with the much enhanced stellar wind produced in the last few years of the pre-supernova stage (Das & Ray 2017), into grid cells within the control volume at the very end of their lifetime.
The lifetime of massive stars is calculated using a scheme similar to that of Gutcke et al. (2021), which involves linearly interpolating the table provided by Portinari et al. (1998) with an assumed metallicity equals to the initial gas metallicity, i.e., $Z=0.02Z_{\odot}$ for most runs. This model has been proven robust with small discrepancies compared to other stellar lifetime models such as Schaller et al. (1992) and Paxton et al. (2010). Note that not all stars with a ZAMS mass above 8 $M_{\odot}$ will end their lives as CCSN. A noticeable fraction of OB stars may collapse directly to form black holes during the core collapse, which is commonly referred to as failed SNe (Woosley & Weaver 1986; MacFadyen & Woosley 1999). Failed SNe often release considerably less energy, typically on the order of $10^{49}$ erg, and result in a much smaller amount of mass return (e.g. Sukhbold et al. 2015). In our simulation, we apply the same mass and energy injection scheme for CCSN to these failed SNe, but with their mass, metal, and energy injection rates inferred from the Sukhbold et al. (2015) model.
Apart from the variable SNe model from Sukhbold et al. (2015), we have also adopted the widely used fixed energy($E_{SN}=10^{51}erg$) SNe model in run fSN, and compare the effect of the two different setups in the following sections.
@@
Radiative cooling and heating play an important role in the evolution of ISM. In our simulation, we use GRACKLE (Smith et al. 2016), an open-source chemistry and cooling library, to solve radiative cooling and heating. GRACKLE provides primordial and dust chemistry calculations, as well as tabulated metallicity-dependent cooling functions with and without self-shielding. Our work adopts a simplified chemistry network that accounts only for gas and metal; a tabulated ionization collisional equilibrium cooling function calculated from CLOUDY is used.
[…]
Figure 2: Face on view of the distribution of star particles (white dots) in the gas disc in the FD simulation, along with the projected gas density. This image shows how star particles are coupled with high-density clumps. The two lower images reveal how super bubbles are centered around clustered particles.
The gas disk of M82 is a highly gas-rich and dust-rich environment, where the optical-thin assumption may fail due to the large opacity in some regions. Previous studies argue that self-shielding can significantly reduce the cooling rate by an order of magnitude in cool to warm regions with relatively low density ($\sim 1\rm{cm^{-3}}$), which is enough for the radiation feedback to become significant to support the clouds from collapsing (Emerick et al. 2019). In our simulation, the self-shielding option provided by GRACKLE is turned on, in which the self-shielding of both HI and HeI from the UV background is included. The detailed derivation of the tabulated self-shielding cooling function in GRACKLE is described in the work of Smith et al. (2016) and Rahmati et al. (2012).
@@ 2.6 Simulations
Our simulations were performed with the state-of-the-art grid-based MHD code Athena++ (Stone et al. 2020). We used the multi-grid self-gravity solver module developed by Kenko Tomida to solve the self-gravity hydrodynamic equations. We also used static mesh refinement (SMR) to refine the cells in the central starburst region to up to 4 pc resolution. The detailed setups of the simulations that we performed are listed in Table 2.
Figure 3: Top: The star formation rate as functions of time in various simulations. Middle: The total mass of stars formed as a function of time. Bottom: The total mass of stars formed in the starburst but still exit, i.e. not explode as SNe, as a function of time.
@@ 3 Starburst in our simulations
M82 has undergone multiple starburst events throughout its history (Divakara Mayya & Carrasco 2009). The majority of stars in the stellar disk are between 300 Myr and 1 Gyr. The recent starburst, which occurred within the past $15\sim 50$ Myr in the nuclear region (within $r<625$ pc), is responsible for the current galactic outflow. Some studies suggest that this recent starburst comprised two distinct events (Rieke et al. 1993; Schreiber et al. 2003), with a major burst peaking around $8\sim 15$ Myr ago and having a peak SFR of $160\,\rm{M_{\odot}}\cdot\rm{yr}^{-1}$ and lasting for only about 1 Myr.
Figure 2 illustrates the star formation process in the FD simulation. At $t=5$ Myr, gravitational instability leads to the formation of numerous small and medium-sized gas clumps, particularly in the nuclear region. Concurrently, a significant number of star particles emerge. As time progresses, these gas clumps grow and spawn more star particles. Clumps merge to form larger clumps, and star particles would form super star clusters. By 15-20 Myr, a substantial number of stars and clusters have formed, leaving only a small volume occupied by dense gas.
@@
Disabling any single stellar feedback mechanism has a minor effect on the SFH. However, disabling all three feedback processes will reduce the SFR by $5\sim 10\%$, which seems somewhat abnormal. Meanwhile, switching to a fixed SNe energy model, which injects roughly twice the energy of the fiducial variable SNe model, moderately reduces the SFR after $t\sim 10$ Myr. The probable reason is that stellar feedback with moderate strength can positively influence star formation. However, excessively strong feedback can suppress star formation activity. We conclude that the overall SFH of the starburst is moderately impacted by the stellar feedback implemented in our simulation. This is likely because, on the starburst timescale of 30 Myr, the initial conditions of the gas disc and the gas return process, which determine the amount of gas can collapse, have a stronger influence.
@@
Figure 3 shows that reducing the gas return fraction from $100\%$ to $50\%$ and $0\%$ significantly lowers both the SFR and SFE, as a substantial fraction of gas becomes locked in stars. Meanwhile, varying the initial metallicity $\rm{Z_{init}}$ notably impacts the SFH. Higher initial metallicity leads to faster cooling, resulting in a more violent starburst during the first 15 Myr, with peak SFRs of $37\,\rm{M_{\odot}\,yr^{-1}}$ in the Z01 simulation and $45\rm{M_{\odot}\,yr^{-1}}$ in Z05. However, this intense early starburst is followed by rapid quenching after $t>15$ Myr, leading to a final star formation efficiency (SFE) and total stellar masses similar to those of the fiducial run (FD). In addition, increasing $\rm{M_{gd}}$ will moderately enhance the SFR at $t>15$ Myr, while a lower $\rm{M_{gd}}$ will suppress star formation activity throughout the simulation.
@@
The middle panel of Figure 3 presents the total stellar mass formed as a function of time. At $t\sim 30$ Myr, the total stellar mass in most simulations ranges from $5.0\times 10^{8}\,\rm{M_{\odot}}$ to $6.0\times 10^{8}\,\rm{M_{\odot}}$, corresponding to a total input of SN energy of $2.3\sim 2.7\times 10^{57}\rm{erg}$ for the variable SN model and $\sim 4.7\times 10^{57}\rm{erg}$ for the fixed SN model. Reducing the gas return fraction to $50\%$ or $0\%$, or adopting a lower gas disk mass of $\rm{M_{gd}}=2.5\times 10^{9}\,\rm{M_{\odot}}$, results in a lower total stellar mass of $3.5\times 10^{8}\,\rm{M_{\odot}}$ to $4.5\times 10^{8}\,\rm{M_{\odot}}$. This implies that approximately $12\%-20\%$ of the initial gas disk mass was converted into stars over 30 Myr. Observational estimates of the total stellar mass formed during the past $\sim 15-30$ Myr range from $2.0\times 10^{8}\,\rm{M_{\odot}}$ to $3.5\times 10^{8}\,\rm{M_{\odot}}$ (Rieke et al. 1993; Schreiber et al. 2003; Strickland & Heckman 2009), which is broadly consistent with our simulations with reduced gas return fractions or lower initial gas disk masses. However, observational estimates of the SFH are subject to uncertainties due to factors such as dust absorption, projection effects, and assumptions about the initial mass function and stellar evolution (e.g.Schreiber et al. 2003).
@@
Another way to compare our results with observational constraints is to examine the total mass of stars formed during the starburst that remain within the nuclear region. The bottom panel of Figure 3 shows the stellar mass within $r<1000$ pc as a function of time in our simulations, suggesting that $\sim 20\%$ of the formed stellar mass has either undergone SN explosions or escaped from the nuclear region. At $t=30$ Myr, the total remaining stellar mass within $r<1000$ pc ranges from $3.0\times 10^{8}M_{\odot}$ to $6.0\times 10^{8}M_{\odot}$, which is below the total mass in the nuclear region (within $\sim 500-600$ pc) estimated by Schreiber et al. (2001).
@@
Figure 4: The surface star formation rate density against the surface gas density in the FD (filled blue circle), fSN (filled yellow circle) and nSN (filled green circle) simulations, with respect to the Kennicutt-Schmidt law (solid blue line), star burst galaxies in Kennicutt 1998 and and young stellar objects in Heiderman et al. 2010.
@@
The properties of star clusters formed during the starburst can significantly influence the launch and development of galactic outflows. Previous studies have employed super star clusters (SSCs) to drive galactic winds in M82-like galaxies (Melioli et al. 2013; Schneider et al. 2020; Schneider & Mao 2024). To investigate the properties of star clusters formed in our simulations, we utilize the friends-of-friends (FoF) algorithm to identify clusters and analyze their spatial distribution, metallicity, mass function (CMF), and other characteristics.
The top panel of Figure 5 presents the cumulative CMF in our simulations. Recent JWST observations byLevy et al. (2024) suggest a power-law CMF for star clusters in M82, $\ y\propto-\alpha\ log\ x$, with a slope of $\alpha=0.9\pm 0.2$. Our simulated CMF exhibits a more bottom-heavy distribution. This discrepancy may be attributed to the imposed mass threshold, $M_{cnv}$, when converting a sink particle to a star particle. The higher-mass end of the CMF, less affected by this threshold, aligns better with the observed power-law slope of $\alpha=0.9$.
While Section 3.1 demonstrates that stellar feedback has a modest impact on the overall SFH, it significantly influences the cumulative CMF. Simulations nRad and nFB exhibit a more top-heavy CMF compared to other simulations, suggesting that stellar feedback effectively disrupts gas accretion in giant star-forming clumps, redistributes gas, and inhibits the formation of massive star clusters. Among the three feedback mechanisms, radiation feedback appears to be the most effective in suppressing the formation of SSCs, reducing the fraction of SSCs with masses exceeding $M_{*}>10^{6}\,\rm{M_{\odot}}$ by $90\%$, and eliminating those with masses above $M_{*}>10^{7}\,\rm{M_{\odot}}$. In contrast, the default variable SNe feedback mode has a more moderate impact on the formation of massive SSCs.
On the other hand, all three stellar feedback mechanisms can suppress the integrated star formation efficiency (SFE) within individual sink particles. The middle panel in 5 across various simulations, which follows a roughly log-normal pattern. The median integrated SFE ranges from $10\%$ to $30\%$, which is generally consistent with previous simulations (Dale et al. 2014; Li et al. 2019; Fukushima & Yajima 2021), but higher than observational estimates of (Chevance et al. 2020). This discrepancy may be partly attributed to our assumption of a $20\%$ star formation efficiency per free-fall time for all sink particles, regardless of their properties.

### v4

@@ 2.4.4 Core-collapse SNe
Conventionally, stars within the zero-age main sequence mass range of $8\sim 30M_{\odot}$ are thought to end their life through the outbreak of core-collapse supernovae (CCSN). Most previous models use a fixed thermal energy injection value of $10^{51}$ erg. However, recent studies indicate that the energy and mass of the ejecta as well as the composition of the ejecta depend on certain parameters, especially the ZAMS mass of the progenitor. In our study, we mainly follow the work of Gutcke et al. (2021). We divide the total mass of each star particle (star cluster) into multiple ZAMS stars according to a Kroupa (2002) IMF. In practice, instead of creating many refined small particles, we store a list of individual virtual stars’ information, including mass and expected lifetime, for each star particle (star cluster). At each time step, we update the age of virtual stars and loop over the lists to locate all the massive stars that have exceeded their expected lifetime and then trigger either a CCSN or failed SN for each such virtual star. The injected energy, ejecta mass, and metal of each SN (failed SN) are then calculated from the Sukhbold et al. (2016) CCSN model accordingly. Stellar wind and UV radiation before the outbreak of CCSN have already cleared the path for SNe Dale et al. (2014). Therefore, we directly inject 100% of SNe’s mass, metal, and energy output, together with the much enhanced stellar wind produced in the last few years of the pre-supernova stage (Das and Ray 2017), into grid cells within the control volume at the very end of their lifetime.
The lifetime of massive stars is calculated using a scheme similar to that of Gutcke et al. (2021), which involves linearly interpolating the table provided by Portinari et al. (1998) with an assumed metallicity equals to the initial gas metallicity, i.e., $Z=0.02Z_{\odot}$ for most runs. This model has been proven robust with small discrepancies compared to other stellar lifetime models such as Schaller et al. (1992) and Paxton et al. (2010). Note that not all stars with a ZAMS mass above 8 $M_{\odot}$ will end their lives as CCSN. A noticeable fraction of OB stars may collapse directly to form black holes during the core collapse, which is commonly referred to as failed SNe (Woosley and Weaver 1986; MacFadyen and Woosley 1999). Failed SNe often release considerably less energy, typically on the order of $10^{49}$ erg, and result in a much smaller amount of mass return (e.g. Sukhbold et al. 2016). In our simulation, we apply the same mass and energy injection scheme for CCSN to these failed SNe, but with their mass, metal, and energy injection rates inferred from the Sukhbold et al. (2016) model.
Apart from the variable SNe model from Sukhbold et al. (2016), we have also adopted the widely used fixed energy ($E_{\rm{SN}}=10^{51}\rm{erg}$) SNe model in run fSN, and compare the effect of the two different setups in the following sections.
@@
Radiative cooling and heating play an important role in the evolution of ISM. In our simulation, we use GRACKLE (Smith et al. 2016), an open-source chemistry and cooling library, to solve radiative cooling and heating. GRACKLE provides primordial and dust chemistry calculations, as well as tabulated metallicity-dependent cooling functions with and without self-shielding. Our simulation use tabulated cooling function calculated from CLOUDY, which account for collisional ionization and ionization from UVB.
The cooling rate of the gas highly depends on the metallicity. For a star-forming gas cloud with constant feedback and metal enrichment, the metallicity can vary from the initial value of 0.02 $Z_{\odot}$ to as high as 20 $Z_{\odot}$, causing the cooling rate to vary by an order of magnitude. This variation in the cooling rate could significantly affect the evolution of molecular clouds, the HII region, and the multiphase outflow. In order to obtain a more accurate result, we use a passive scalar as the tracer for metal elements in the ISM. Moreover, we include the SNe metal enrichment based on the isotope yield provided in the work of Sukhbold et al. (2016). Therefore, we can trace the detailed evolution of the metal elements in the ISM and calculate an accurate cooling function accordingly.
Figure 2: Face-on view of the gas disk in the FD simulation, showing the projected gas density overlaid with star particles (white dots). This picture shows star particles form predominantly within high-density clumps. The bottom panels illustrate how superbubbles develop around clustered star particles. In these panels, the large number of star particles partly obscures the underlying high-density clumps.
The gas disk of M82 is a highly gas-rich and dust-rich environment, where the optical-thin assumption may fail due to the large opacity in some regions. Previous studies argue that self-shielding can significantly reduce the cooling rate by an order of magnitude in cool to warm regions with relatively low density ($\sim 1\rm{cm^{-3}}$), which is enough for the radiation feedback to become significant to support the clouds from collapsing (Emerick et al. 2019). In our simulation, the self-shielding option provided by GRACKLE is turned on, in which the self-shielding of both HI and HeI from the UV background (UVB) is included. The Haardt and Madau (2012) cosmic UV background model is adopted to mimic the UVB. The detailed derivation of the tabulated self-shielding cooling function in GRACKLE is described in the work of Smith et al. (2016) and Rahmati et al. (2012).
@@ 2.6 Simulations
Our simulations were performed with the state-of-the-art grid-based MHD code Athena++ (Stone et al. 2020). In our simulation, we use the Harten–Lax–van Leer–Contact (HLLC) scheme as the Riemann solver. For self-gravity, we used the multi-grid self-gravity solver module developed by Tomida and Stone (2023) to solve the self-gravity hydrodynamic equations. At present, our simulations do not incorporate magnetic fields. The simulations were run in a cubic box with a side length of 4 kpc and an outflow boundary condition. Static mesh refinement (SMR) is used to refine the cells in the central starburst region to up to 4 pc resolution. The detailed description of the SMR scheme is given in Appendix B. Note that the gas disk is not truncated in our simulations. Although not shown in the manuscript, we performed additional test runs and confirmed that this choice does not affect our results. The detailed setups of the simulations that we performed are listed in Table 2.
Figure 3: Top: The star formation rate as a function of time in various simulations. Middle: The cumulative mass of stars formed since the beginning of the simulation. Bottom: The remaining stellar mass within $r<1000$ pc that has yet undergone SNe as a function of time.
@@ 3 Starburst in our simulations
M82 has undergone multiple starburst events throughout its history (Divakara Mayya and Carrasco 2009). The majority of stars in the stellar disk are between 300 Myr and 1 Gyr. The recent starburst, which occurred within the past $15\sim 50$ Myr in the nuclear region (within $r<625$ pc), is responsible for the current galactic outflow. Some studies suggest that this recent starburst comprised two distinct events (Rieke et al. 1993; Schreiber et al. 2003), with a major burst peaking around $8\sim 15$ Myr ago and having a peak SFR of $160\,\rm{M_{\odot}}\cdot\rm{yr}^{-1}$ and lasting for only about 1 Myr.
Figure 2 illustrates the star formation process in the FD simulation. At $t=5$ Myr, gravitational instability leads to the formation of numerous small and medium-sized gas clumps, particularly in the nuclear region. Concurrently, a significant number of star particles emerge. As time progresses, these gas clumps grow and spawn more star particles. Clumps merge to form larger clumps, and star particles would form super star clusters. By 15-20 Myr, a substantial number of stars and clusters have formed, leaving only a small volume occupied by dense gas. We find that most of the stars form within $r<1.0$ kpc, while hardly any stars form beyond $r>1.2$ kpc due to the low gas density in that region.
@@
Disabling any single stellar feedback mechanism has a minor effect on the SFH. However, disabling all three feedback processes will reduce the SFR by $5\sim 10\%$, which seems somewhat abnormal. Meanwhile, switching to a fixed SNe energy model, which injects roughly twice the energy of the fiducial variable SNe model, moderately reduces the SFR after $t\sim 10$ Myr. The probable reason is that stellar feedback with moderate strength can positively influence star formation. However, excessively strong feedback can suppress star formation activity. We conclude that the stellar feedback has a minor impact on the overall SFH of the starburst. This is likely because, on the starburst timescale of 30 Myr, the amount of gas that can collapse and how quickly the gas can collapse is mostly determined by the initial conditions of the gas disc. Additionally, the gas return process implemented in most simulations allows particles to release the accreted gas mass back to the ISM, effectively sustaining star formation activity at later time.
@@
Figure 3 shows that reducing the gas return fraction from $100\%$ to $50\%$ and $0\%$ significantly lowers both the SFR and overall SFE within 30 Myr, as a substantial fraction of gas becomes locked in sink/star particles. Meanwhile, varying the initial metallicity $\rm{Z_{init}}$ notably impacts the SFH. Higher initial metallicity leads to faster cooling, resulting in a more violent starburst during the first 15 Myr, with peak SFRs of $37\,\rm{M_{\odot}\,yr^{-1}}$ in the Z01 simulation and $45\,\rm{M_{\odot}\,yr^{-1}}$ in Z05. However, this intense early starburst is followed by rapid quenching after $t>15$ Myr, leading to a final SFE and total stellar masses similar to those of the fiducial run (FD). In addition, increasing $\rm{M_{gd}}$ will moderately enhance the SFR at $t>15$ Myr, while a lower $\rm{M_{gd}}$ will suppress star formation activity throughout the simulation.
@@
The middle panel of Figure 3 presents the total stellar mass formed as a function of time. At $t\sim 30$ Myr, the total stellar mass in most simulations ranges from $5.0\times 10^{8}\,\rm{M_{\odot}}$ to $6.0\times 10^{8}\,\rm{M_{\odot}}$, corresponding to a total input of SN energy of $2.3\sim 2.7\times 10^{57}\rm{erg}$ for the variable SN model and $\sim 4.7\times 10^{57}\rm{erg}$ for the fixed SN model. Reducing the gas return fraction to $50\%$ or $0\%$, or adopting a lower gas disk mass of $\rm{M_{gd}}=2.5\times 10^{9}\,\rm{M_{\odot}}$, results in a lower total stellar mass of $3.5\times 10^{8}\,\rm{M_{\odot}}$ to $4.5\times 10^{8}\,\rm{M_{\odot}}$. This implies that approximately $12\%-20\%$ of the initial gas disk mass was converted into stars over 30 Myr. Observational estimates of the total stellar mass formed during the past $\sim 15-30$ Myr range from $2.0\times 10^{8}\,\rm{M_{\odot}}$ to $3.5\times 10^{8}\,\rm{M_{\odot}}$ (Rieke et al. 1993; Schreiber et al. 2003; Strickland and Heckman 2009), which is broadly consistent with our simulations with reduced gas return fractions or lower initial gas disk masses. However, observational estimates of the SFH are subject to uncertainties due to factors such as dust absorption, projection effects, and assumptions about the initial mass function and stellar evolution (e.g.Schreiber et al. 2003).
@@
Another way to compare our results with observational constraints is to examine the total mass of stars formed during the starburst that remain within the nuclear region. The bottom panel of Figure 3 shows the stellar mass within $r<1000$ pc as a function of time in our simulations, suggesting that $\sim 20\%$ of the formed stellar mass has either undergone SN explosions or escaped from the nuclear region. At $t=30$ Myr, the total remaining stellar mass within $r<1000$ pc ranges from $3.0\times 10^{8}M_{\odot}$ to $6.0\times 10^{8}M_{\odot}$, which is below the stellar mass inferred from observations in the nuclear region (within $\sim 500-600$ pc), as reported by Schreiber et al. (2001).
@@
Figure 4: The surface star formation rate density against the surface gas density in the FD (filled blue circle), fSN (filled yellow circle) and nSN (filled green circle) simulations, compared to the to the Kennicutt-Schmidt law (solid blue line), starburst galaxies and spiral galaxies in Kennicutt 1998. For clarity, the fSN and nSN data points have been shifted upward by 0.5 and 1.0 dex, respectively.
@@
The properties of star clusters formed during the starburst can significantly influence the launch and development of galactic outflows. Previous studies have employed super star clusters (SSCs) to drive galactic winds in M82-like galaxies (Melioli et al. 2013; Schneider et al. 2020; Schneider and Mao 2024). To investigate the properties of star clusters formed in our simulations, we utilize the friends-of-friends (FoF) algorithm to identify clusters and analyze their spatial distribution, metallicity, mass function (CMF), and other characteristics. Considering the limited spatial resolution, here we used a linking length $l_{\rm{fof}}=8.0\,\rm{pc}$.
The top panel of Figure 5 presents the cumulative CMF in our simulations. Recent JWST observations by Levy et al. (2024) suggest a power-law CMF for star clusters in M82, $y\propto\alpha\rm{log}(x)$, with a slope of $\alpha=0.9\pm 0.2$. Our simulated CMF exhibits a more bottom-heavy distribution. This discrepancy may be attributed to the imposed mass threshold, $M_{cnv}$, when converting a sink particle to a star particle. The higher-mass end of the CMF, less affected by this threshold, aligns better with the observed power-law slope of $\alpha=0.9$.
While Section 3.1 demonstrates that stellar feedback has a limited impact on the overall SFH, it significantly influences the cumulative CMF. Simulations nRad and nFB exhibit a more top-heavy CMF compared to other simulations, suggesting that stellar feedback effectively disrupts gas accretion in giant star-forming clumps, redistributes gas, and inhibits the formation of massive star clusters. Among the three feedback mechanisms, radiation feedback appears to be the most effective in suppressing the formation of SSCs, reducing the fraction of SSCs with masses exceeding $M_{*}>10^{6}\,\rm{M_{\odot}}$ by $90\%$, and eliminating those with masses above $M_{*}>10^{7}\,\rm{M_{\odot}}$. In contrast, the default variable SNe feedback mode has a more moderate impact on the formation of massive SSCs.
On the other hand, all three stellar feedback mechanisms can suppress the integrated star formation efficiency within individual sink particles. The middle panel in 5 shows the integrated SFE of sink particles across various simulations, which follows a roughly log-normal pattern. The median integrated SFE ranges from $10\%$ to $30\%$, which is generally consistent with previous simulations (Dale et al. 2014; Li et al. 2019; Fukushima and Yajima 2021), but higher than observational estimates of (Chevance et al. 2020). This discrepancy may be partly attributed to our assumption of a $20\%$ star formation efficiency per free-fall time for all sink particles, regardless of their properties.